A sample of hard water contains 122 ppm of HCO3- ions. What is the minimum weight of CaO required to remove ions completely from 1 kg of such water sample?
 
  • 56 mg
  • 112 mg
  • 168 mg
  • 244 mg

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Temporary hardness is due to the HCOɵ3 of Ca2+ and Mg2+ and is removed by either boiling or by treating with CaO or Ca(OH)2 (Clark's method) 
Ca(HCO3)2+CaO => 2CaCO3 + H2O ..(i) 
122p± of HCO3^-1 =  122gHCO3 in 10^6mL or 10^3L 
122ml HCO3 in 1.0L 
122/61=2mmol in 1.0L 
2/2 n-factor =2 
=1m mol of Ca(HCO3)2 in 1.0L 
According to equation (i), 
1m mol of Ca(HCO3)2≡1m mol of CaO in 1.0L 
1×56=56 mg of CaO 
56ofCaO in 1.0L 
56 mg of CaO in 1kgofH2O 
(1LofH2O = 1kg of H2O ∵dH2O=1)